# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v1

@@ 5 Additional Analysis of LASeR
Robustness to Noisy Rewards.
To examine the robustness of our method in the presence of noisy or irrelevant rewards, we conduct the following analysis using Llama-3-8B on GSM8K. We add varying amounts of Gaussian noise $\sigma$ to the rewards generated by RMs sampled from the distribution $\mathcal{N}(0,\sigma I)$, where $I$ is the identity matrix, to simulate noisy rewards when using RMs in out-of-distribution settings. In addition to LASeR using the LinUCB algorithm (c.f. Sec. 3.2),
we also use Exp3 (Auer et al., 2002) designed for adversarial bandit settings. In Fig. 3, we find that even as the degree of noise in RM scores increases (from $\sigma\!=\!$ 0.1 to 0.4), LASeR’s selection strategy continues to perform robustly, mitigating the effects of noise compared to the sequential baseline. Specifically, LASeR has an average performance drop of only $0.55\%$ accuracy at a noise level of $\sigma\!=\!0.3$, whereas the sequential baseline suffers a $1.6\%$ accuracy drop (3 times as much) under the same conditions. Furthermore, LASeR using Exp3, the most noise-robust method, maintains consistent performance, with only a $0.26\%$ accuracy drop.
Figure 4: Training efficiency of LASeR (LinUCB, and Exp3) vs. sequential selection and RM ensemble on GSM8K.
@@
As we noted in Sec. 4.1, standard multi-reward baselines such as sequential and random RM selection are slow to converge. We now concretely show the accuracy-training time tradeoff in Fig. 4 by comparing the GSM8K performance of training with LASeR (with LinUCB and Exp3) with sequential RM selection and the offline RM ensemble, along with the corresponding wall clock training time. [^5] Wall clock time is measured as the training time of a model (hours), keeping compute resource consistent. We find that sequentially optimizing over each RM performs the worst in terms of training time (3×3$\times$3 × of LASeR) while offline RM ensemble has the worst accuracy (and takes 2×2$\times$2 × the training time of LASeR). Among the MAB algorithms, LinUCB achieves the best balance. Exp3, while slower than LinUCB, reaches the highest accuracy, demonstrating its strength in effectively managing exploration across RMs.
Figure 5: Agreement in preference rankings between RMs on output pairs generated by Llama-3-8B on MMLU (left) and WildChat (right).
Presence of Conflicting Signals among RMs.
In Sec. 4.2, we find that LASeR consistently outperforms other multi-reward baselines across a wide variety of tasks. We attribute some of these performance gains to the inability of the multi-reward baseline to handle conflicting signals, resulting in subpar training data from multiple RMs. To study this, we sample pairs of outputs generated by Llama-3-8B on MMLU as well as WildChat and evaluate the consistency of response preferences measured by multiple RMs. Since pair-wise preferences are binary, we compute F1 to measure consistency with one RM’s preferences serving as the reference. Fig. 5 (left on MMLU) reveals that Qwen and Zephyr have the highest agreement rate at $0.77$, while Qwen’s agreement with Eurus and Olmo is lower at $0.58$ and $0.43$, respectively. Zephyr also shows low agreement with Olmo at $0.57$. This is expected as Qwen and Zephyr are the top-performing models in reasoning according to RewardBench, while Olmo ranks the lowest in reasoning ability among the four models. We observe similar trends in agreement across RMs on WildChat (albeit with different agreement scores), which contains user queries asked LLMs in the wild; see Fig. 5 (right). It appears that for more heterogeneous datasets with more categories, the level of disagreement among RMs (or conflict) increases. This also highlights the benefit of LASeR over multi-RM baselines that do not adequately address conflicts in RMs.
LASeR’s selected RM adjusts to the Query.
Generalization ability of LASeR.
While recent work focuses on building RMs that reflect preferences across domains, a large body of prior work developed a suite of evaluation metrics catered to specific domains such as reasoning (Golovneva et al., 2022; Prasad et al., 2023). In Table 5 (Appendix C), we show that Llama-3-8B trained using LASeR, to adaptively select relevant evaluation metrics, outperforms baselines by $1.62\%$ on average and can effectively filter underperforming metrics without degrading performance (c.f. Fig. 8). Furthermore, from Table 6, we observe that on reasoning datasets LASeR outperforms sequential RM optimization under four different choices of loss functions: NLL, DPO (Rafailov et al., 2024b), DPO + NLL (Pang et al., 2024), and KTO (Ethayarajh et al., 2024). Finally, in Table 7, we find that models trained with LASeR also exhibit the highest generalization to out-of-distribution settings such as on CommonsenseQA (Talmor et al., 2018) and MATH (Hendrycks et al., 2021c), reiterating the broad generalizability of LASeR.
6 Discussion and Conclusion
LASeR with different “kinds” of RMs.
In Sec. 4.2, we show that LASeR can choose from a set of candidate RMs, and our analysis in Fig. 3 highlights the fact that LASeR is robust to noisy RMs. Furthermore, in Appendix C, we show that LASeR can also be used with metric-based rewards (Golovneva et al., 2022). These experiments reflect a conceptual split between the generator (the LLM) and the scorer (the RM or metric). Thus, LASeR is applicable to other settings that follow this paradigm, e.g., using an LLM-as-a-judge (Zheng et al., 2023), where LASeR could be used to choose between different judge models or prompts or different combinations of RMs and metrics. We leave adding such models into LASeR to future work.
Quality of RMs used with LASeR.
Methods that rely on RMs for scoring generally assume that these RMs have a strong correlation with human judgments. LASeR tempers this assumption in a number of ways: first, by ensembling multiple RMs, LASeR weakens the effect of noisy RMs; this can be seen in Fig. 3, where LASeR mitigates the negative impact of a noisy RM even as the level of noise is increased. Moreover, the fact that LASeR can select RMs at an instance-level means that there need not be a single RM that always correlates well across all instances. However, LASeR does require at least one RM to be positively correlated with human judgments on each instance. If this assumption is not met (i.e., all RMs are poorly-correlated across all instances), then optimizing for the RMs will yield poor results. Note that this holds true for any method optimizing for RMs. Because LASeR selects from multiple RMs, its contributions are complementary to developments in RMs, which can easily be integrated into LASeR, as well as improvements to preference optimization loss functions (see Appendix C). Such RM improvements are likely to be necessary as LLMs are deployed in domains that are out of scope for existing systems and domains with heterogeneous requirements (e.g., our generation domains in Sec. 4.2). In these cases, there will be no single “perfect” existing RM, and successful solutions will likely involve mixing multiple RMs. A core benefit of LASeR is its the ability to automatically filter RMs; in Fig. 6 we see that utilization differs across domains. This allows users to avoid expensive experimentation with subsets of RMs: they can simply offload this task to LASeR, which will automatically select the more useful RM(s).
Conclusion.
@@
We present LASeR, an adaptive method for selecting RMs and iteratively training LLMs using multiple RMs. We formulate RM selection as a contextual multi-armed bandit problem, learning to select the RM that most improves the LLM conditioned on the given input or query. We test LASeR across diverse settings, demonstrating its utility on reasoning tasks, open-ended instruction-following tasks, as well as long-context generation, where we use the selected RM to rerank outputs without training the LLM. Across domains, we show that LASeR consistently results in superior performance, whereas multi-RM baselines that select RMs using random or fixed strategies, or ensemble multiple RMs uniformly, have lower and more variable performance. In our analysis, we show that LASeR is robust to noisy RMs, and flexibly uses different RMs depending on the domain, and generalizes to multiple settings. Moreover, by selecting one RM at a time, LASeR provides the best of both worlds: consistently outperforming all baselines while still maintaining efficiency by only optimizing for one model at a time. Taken together, our results indicate that LASeR is a powerful tool for improving LLMs, providing us with the robustness and performance improvements associated with training against multiple RMs without incurring the downsides of having to select only one fixed RM or having to handle conflicting reward signals.
@@ Appendix A Experiments
A.1 Experimental Settings
Training setup. For training with LoRA, we set the rank to 16 and alpha to 32. We fine-tune the model for 10 iterations using a learning rate of $65e\!-\!$ 6 and a batch size of 16. Following Pang et al. (2024), we generate $P=10$ pairs per problem for training with our loss in Sec. 3.1. For all experiments, we trained each method to converge. The number of iterations for each method was selected based on the observed convergence, with a performance metric threshold of $0.1$ across training batches on the dev set. Our experiments are run on 4 RTX A6000 with 48G memory each.
@@ A.1 Experimental Settings
RewardBench. Following Lambert et al. (2024), rewards are computed with no reference model and only used the log-likelihood of the reward model. For instance, given a reward model $\pi_{R^{\star}}$, the reward for an input $x_{i}$ and response $y_{i}$ is calculated as: $\log\pi_{R^{\star}}(y_{i}\mid x_{i})$. There is no need for normalization since we use this log-likelihood to rank the responses. Specifically, on GSM8K, we computed the average and standard deviation of the rewards for the chosen responses as $-4.2965$ and $1.4156$, respectively, and for the rejected responses as $-5.9861$ and $1.6546$, respectively.
@@
Datasets. For StrategyQA, GSM8K, and MMLU, we divided each dataset into training and test sets. The model is fine-tuned on the training set and dev set, then evaluate on the test set. For WildChat, the dataset was split into a 70/10/20 ratio for training, development, and testing. Following Zhao et al. (2024), prompt categorization is done using a small off-the-shelf classifier. [^6] Link: https://huggingface.co/valpy/prompt-classification For LongBench, we subsample 5K examples for three tasks: multi-document QA, summarization, and few-shot learning. Each category was split into a 70/10/20 ratio, and the bandit model was trained and validated on the training and development sets and then tested on the test set. We report the detailed number of instances for train, development, and test sets in Table 3.
Prompts. We present a collection of prompt templates designed for each dataset as follows:
@@ A.2 Details of Bandit Algorithms
MAB reward normalization. To maintain a consistent scale and magnitude MAB rewards across training, we apply scaled rewards based on the quantiles of the reward history, following the method outlined by Graves et al. (2017). Let $L=\{-\mathcal{L}^{m}(1),\dots,-\mathcal{L}^{m}(t-1)\}$ represent the unscaled reward history up to time step $t$. This history’s lower and upper quantiles are denoted as $q_{t}^{lo}$ and $q_{t}^{hi}$, respectively. We set $q_{t}^{\text{lo}}$ and $q_{t}^{\text{hi}}$ to be $20^{th}$ and $80^{th}$ quantiles. The scaled reward, ^$-\hat{\mathcal{L}}^{m}(t)$, is then defined as:
@@ Appendix B Additional empirical results
LASeR adaptively selects from multiple RMs on LongBench. On LongBench (Fig. 7), we observe distinct utilization patterns for the QA tasks vs. summarization and few-shot learning. QA tasks exhibit nearly equal utilization of the top-2 RMs on RewardBench (Zephyr-7B-Alpha and Qwen1.5-7B-Chat in decreasing order), with the utilization of the Qwen RM even exceeds that of Zephyr RM for multi-document QA. In contrast, on summarization and few-shot learning the top RM (Zephyr) is far more preferred by LASeR with margins of $59\%$ and $31\%$ over the second-best RM and the least performant RM being utilized less that $3\%$ of the times.
Detailed results for each RM. Here, we provide detailed reasoning results for each chosen RM where we use a single RM during training (c.f. Sec. 3.1) in Table 4. These results demonstrate that Zephyr-7b-alpha outperforms other RMs on StrategyQA and MMLU, except for GSM8K, where Qwen1.5-7B-Chat has the best performance.

### v2

@@ 5 Additional Analysis of LASeR
Presence of Conflicting Signals among RMs.   In Sec. 4.2, we find that LASeR consistently outperforms other RM ensemble baselines across a wide variety of tasks. We attribute some of these performance gains to the inability of the multi-RM baseline to handle
Figure 3: Agreement in preference rankings between RMs on MMLU (left) and WildChat (right).
conflicting signals, resulting in subpar training data from multiple RMs. To study this, we sample pairs of outputs generated by Llama-3-8B on MMLU as well as WildChat and evaluate the consistency of response preferences measured by multiple RMs. Since pair-wise preferences are binary, we compute F1 to measure consistency with one RM’s preferences serving as the reference. Fig. 3 (left on MMLU) reveals that Qwen and Zephyr have the highest agreement rate at $0.77$, while Qwen’s agreement with Eurus and Olmo is lower at $0.58$ and $0.43$, respectively. This is expected as Qwen and Zephyr are the top-performing models in reasoning according to RewardBench, while Olmo ranks the lowest in reasoning ability among the four models. We observe similar trends in agreement across RMs on WildChat (albeit with different agreement scores), which contains user queries asked LLMs in the wild; see Fig. 3 (right). It appears that for more heterogeneous datasets with more categories, the level of disagreement among RMs (or conflict) increases. This also highlights LASeR’s advantages over multi-RM baselines that do not address conflicts in RMs and may explain why choosing one RM in LASeR and the best RM baseline outperforms multi-RM ensembles.
Figure 4: Training efficiency of LASeR vs. different baselines on GSM8K.
@@
We show the accuracy-training time tradeoff in Fig. 4 by comparing the GSM8K performance of training with LASeR and different baselines, along with the corresponding wall clock training time. Wall clock time is measured as the training time of a model (hours), keeping compute resource consistent. We find that sequentially optimizing over each RM performs the worst in terms of training time (3×3$\times$3 × of LASeR) while RM score ensemble has the worst accuracy (and takes 2×2$\times$2 × the training time of LASeR). Moreover, LASeR outperforms all other baselines in terms of accuracy while maintaining the lowest training time, being more than twice as fast as the second-best baseline. LASeR’s efficiency comes from fast convergence compared to the sequential selection baseline and avoiding the overhead of loading and evaluating multiple RMs per step, as required in RM ensemble methods (Online Ensemble, Score Ensemble, and RM Agreement Ensemble). Instead, it selects and uses only one RM per mini-batch, saving GPU memory and compute. Ensemble baselines require scoring all candidate responses with all RMs, whereas LASeR scores responses with only the selected RM, reducing overhead. Lastly, we note that LASeR can be run in an offline data creation mode, where RM selection and scoring are done once to generate preference data, unlike online ensemble methods that require multiple RM passes per step.
6 Conclusion
@@
We present LASeR, an adaptive method for selecting RMs and iteratively training LLMs using multiple RMs. We formulate the problem as a contextual multi-armed bandit problem, learning to select the RM that most improves the LLM conditioned on the given input or query. We test LASeR across diverse settings, showing its utility on reasoning tasks, instruction-following tasks, and long-context generation. Across domains, we show that LASeR consistently results in superior performance, whereas multi-RM baselines that select RMs using random or fixed strategies or ensemble multiple RMs uniformly have lower and more variable performance. In our analysis, we show that LASeR is robust to noisy RMs, and flexibly uses different RMs depending on the domain, and generalizes to multiple settings. Lastly, by selecting one RM at a time, LASeR provides the best of both worlds: consistently outperforming all baselines while still maintaining efficiency by only optimizing for one model at a time.
Limitations and Broader Impacts
Limitations.
While LASeR demonstrates strong empirical performance and generalization across tasks, policy models, and RMs, it still has some limitations. First, like all work using RMs to improve LLMs, it relies on the availability of a high-quality pool of RMs, and downstream performance may degrade if the pool contains only weak or biased models. Additionally, although the bandit-based approach mitigates reliance on a single RM, it does not explicitly address issues of safety, fairness, or bias in RM selection and consider potential trade-offs between safety and task effectiveness when choosing an RM. Finally, besides the MAB formulation, other methods such as reinforcement learning-based RM selection might be applicable. This could allow for more flexible RM selection policies, but would introduce substantial optimization and stability challenges, especially when dealing with non-stationary and noisy RM feedback. We consider these directions as promising for future research.
Broader Impact.
LLMs have been shown to reflect stereotypes, biases, and other negative traits contained in their pretraining data (Weidinger et al., 2021). Consequently, finetuned LLMs (including those trained with LASeR) may also exhibit such undesirable traits in their generations during inference or training and exhibit the same potential for misuse as any other finetuned model. While prior work has made some headway in detecting such harmful content generated by LLMs (Inan et al., 2023), considerable research effort is needed in mitigating bias in LLMs. Conceptually, classifiers that detect risky, harmful, or biased content in the text can also be used as an additional RM in LASeR’s training to reinforce avoiding bias via preference optimization. However, we do not study this in our work and leave it to future work to explore these directions.
@@ Appendix A Experiments
A.1 Experimental Setting
Training Setup.
We provide the training details for our method and baselines as follows:
  • LoRA: For training with LoRA, we set the rank to 16 and alpha to 32.
  • Preference pairs construction: Following Pang et al. [2024], we generate $P=10$ pairs per query for training with our loss in Sec. 3.1.
  • Training iterations: For all experiments, we trained each method to converge. The number of iterations is selected based on the observed convergence, with a performance metric threshold of $0.1$ across training batches on the dev set. In particular, LASeR, “Best RM”, “Avg. RM”, “Classifier RM”, and RM ensemble baselines were trained for 10 iterations. For both the sequential and random RM selection, we found LLM training took longer to converge, and consequently, the model was trained for 25 iterations.
  • Batch size: We fine-tune the model using a learning rate of $65e\!-\!$ 6 and a batch size of 16. Noted that we experimented with different batch sizes to evaluate their impact. Using a batch size of 1 yielded comparable performance to a batch size of 16 but was significantly less efficient in training the LLM due to the increased computational overhead. Based on this, we opted to use a batch size of 16 for a better trade-off between performance and efficiency. For datasets like Wildchat, which contain clearly defined and diverse categories, we structure the batches such that each batch consists of data belonging to a single category. This setup minimizes the risk of mismatches between the RM and the batch data, as each RM is evaluated on its most relevant data category. For reasoning datasets where such predefined categories do not exist, shuffling the data during training ensures diverse data and a good training signal for LLM within each batch.
  • Resources: The LinUCB algorithm has a total of 1.6M learnable parameters (including matrix $A$ and bias vector $b$). Regarding the computation of MAB parameters, empirically, we find that for query embedding with dimension $4096$ and using 4 reward models, inverse computation needs to be performed for each RM once per batch (and can be cached to be reused later), which only adds a latency of $1.27$ seconds. In comparison, a forward and backward pass of the LLM takes $36.39$ seconds. Our experiments are run on 4 RTX A6000 with 48G memory each.
@@ A.1 Experimental Setting
RewardBench. Following Lambert et al. [2024], rewards are computed with no reference model and only use the log-likelihood of the reward model. For instance, given a reward model $\pi_{R^{\star}}$, the reward for an input $x_{i}$ and response $y_{i}$ is calculated as: $\log\pi_{R^{\star}}(y_{i}\mid x_{i})$. There is no need for normalization since we use this log-likelihood to rank the responses.
Figure 5: Reward difference distribution of chosen and rejected responses of Llama-3-8B on RewardBench.
To test the hypothesis that the log probabilities of the chosen and rejected responses from the policy model are generally close in value, with the chosen response tending to have a slightly higher log probability, we conduct an experiment using Llama-3-8B and RewardBench dataset. For each instance, we calculated the reward difference by subtracting the log probability (reward) of the chosen response from that of the rejected response. Formally, given a policy model $\pi_{0}$, a prompt $x_{i}$ and a pair of chosen response $y_{i}^{w}$ and rejected response $y_{i}^{l}$ from RewardBench data, we calculate the reward difference as $\log\pi_{0}(y_{i}^{w}\mid x_{i})-\log\pi_{0}(y_{i}^{l}\mid x_{i})$. As shown in Fig. 5, the distribution is centered near zero, indicating that the values are often close. However, the mass is skewed slightly to the right, suggesting that the chosen response tends to have a higher log probability than the rejected one in the majority of cases. This suggests the policy model generally has a reasonable ranking, but benefits from DPO’s maximization of the margin.
Details of RMs. We provide details for each chosen RM:
  • Zephyr-7B-Alpha: is a fine-tuned version of Mistral-7B model that was trained on on UltraChat [Ding et al., 2023] and UltraFeedback [Cui et al., 2023] using DPO.
  • Qwen1.5-7B-Chat: is pretrained with human-style conversation data inspired by Ouyang et al. [2022] along with questions, instructions, and answers in natural language, and post-trained with both SFT and DPO using diverse prompts [Lu et al., 2023].
  • Eurus-7B-KTO: is a fine-tuned version of Eurus-7B-SFT model using KTO loss on UltraInteract [Yuan et al., 2024a] and UltraFeedback [Cui et al., 2023].
  • OLMo-7B-Instruct: is the instruct version of OLMo-7B base model and was fine-tuned using UltraFeedback [Cui et al., 2023].
Extracting Embeddings for a Query Using $\pi_{m}$. To extract embeddings for a query using $\pi_{m}$, we first process the input query through the policy model $\pi_{m}$. We use the embedding of the last token in the query as the representation for the query. The embedding is then used as input to the subsequent bandit algorithm.
@@
Datasets. For StrategyQA, GSM8K, and MMLU, we divided each dataset into training and test sets. The model is fine-tuned on the training set and dev set and then evaluated on the test set. For WildChat, the dataset was split into a 70/10/20 ratio for training, development, and testing. Following Zhao et al. [2024], prompt categorization is done using a small off-the-shelf classifier. [^5] Link: https://huggingface.co/valpy/prompt-classification For LongBench, we subsample 5K examples for three tasks: multi-document QA, summarization, and few-shot learning. Each category was split into a 70/10/20 ratio, and the bandit model was trained and validated on the training and development sets and then tested on the test set. We report the detailed number of instances for train, development, and test sets in Table 3.
Baselines. Here we provide more details for baselines:
  • Classifier Selection. We add an additional baseline that uses the RewardBench data to train a classifier that maps queries to an RM $\mathcal{C}:\mathbb{R}^{d}\rightarrow\mathcal{R}$, where $\mathcal{R}={R_{1},R_{2},\dots,R_{K}}$ is the set of RMs. Specifically, to construct a dataset for training $\mathcal{C}$, we take each query in the RewardBench data along with its corresponding chosen and rejected responses. The RewardBench dataset contains a total of 2985 examples across several categories, including chat, safety, and reasoning. The dataset is split into an 80/20 ratio for training/development sets, then the classifier is trained on the training set and validated on the development set. We use each RM to score these responses. The RM that assigns the correct score with the highest difference between the chosen and rejected response is selected to label the RM for that query. After training $C$, we use this classifier to select the RM used for training the LLM in our pipeline. In the experiments, we use a three-layer MLP with hidden dimensions of 2048 and 1024 and an output dimension of 4 (number of RMs), with ReLU activation in each layer.
  • RM Ensembles. While the ensemble methods generate scores from multiple RMs in a single iteration for a fixed set of responses sampled at the start of the iteration, we still generate new responses at each training iteration as part of the overall learning process. This ensures that the training dynamically incorporates updated responses from the LLM.
Licenses for Models and Datasets. Below is the license for models and datasets used in this paper:
  • LLMs: Llama-3-8B (Llama 3 Community License), Mistral-7B (Apache 2.0 License), Qwen2.5-32B (Apache 2.0 License).
  • Datasets: StrategyQA, GSM8K, MMLU (MIT License), WildChat (ODC-BY License), LongBench (Apache 2.0 License).
Additional Results and Analysis. In Appendix B, we present additional analyses on the statistical stability of LASeR and baselines across multiple runs (Table 4). We show that the RM selected by LASeR adapts to input queries (Fig. 6) and aligns well with RewardBench ground truth rankings (Fig. 8). We also show that MAB rewards converge over training (Fig. 9), and that the learned MAB policy can be reused to improve cold-start training (Table 5). Detailed comparisons against individual RMs used in isolation (Tables 6 and 7) further confirm that LASeR consistently outperforms all single-RM baselines. In addition, we show that LASeR maintains strong performance across different sets of RMs (Table 8).
In Appendix C, we evaluate the generalization of LASeR across multiple dimensions: different bandit algorithms (Table 9), base models and RMs (Tables 10 and 12), number of RMs (Fig. 11), training loss functions (Table 13), and out-of-distribution tasks (Table 14). We also show that LASeR can successfully select evaluation metrics tailored to domain-specific reasoning tasks [Golovneva et al., 2022, Prasad et al., 2023] (Table 15). Finally, LASeR demonstrates robustness to both noisy RMs and underperforming reward signals (Figs. 12 and 13).
@@ A.2 Details of Bandit Algorithms
MAB reward normalization. To maintain a consistent scale and magnitude of MAB rewards across training, we apply scaled rewards based on the quantiles of the reward history, following the method outlined by Graves et al. [2017]. Let $L=\{-\mathcal{L}^{m}(1),\dots,-\mathcal{L}^{m}(t-1)\}$ represent the unscaled reward history up to time step $t$. This history’s lower and upper quantiles are denoted as $q_{t}^{lo}$ and $q_{t}^{hi}$, respectively. We set $q_{t}^{\text{lo}}$ and $q_{t}^{\text{hi}}$ to be $20^{th}$ and $80^{th}$ quantiles. The scaled reward, ^$-\hat{\mathcal{L}}^{m}(t)$, becomes:
@@
We chose to use the cumulative MAB reward because it provides a more comprehensive measure of how well the algorithm performs over time. In our framework, cumulative MAB reward reflects the performance of the model across batches and iterations, capturing the long-term impact of both exploration and exploitation decisions. Algorithms like LinUCB are designed to optimize cumulative rewards, as they ensure a balanced trade-off between exploration (gathering information about less-tested RMs) and exploitation (using the best-known RM).
@@ Appendix B Additional Empirical Results
Analysis on Statistical Significance of LASeR.
While our main results are based on single runs, we note that this is standard practice in iterative preference optimization setups, including recent works such as Pang et al. [2024], where large-scale LLM training limits the feasibility of multi-seed replication across all benchmarks. Nonetheless, to provide some insights regarding the statistical significance of LASeR, we report the mean and standard deviation over 5 seeds on reasoning tasks. We compare LASeR against the Best RM and RM Agreement Ensemble (identical random seeds to ensure fair comparison), which are the most competitive across our benchmarks. Table 4 shows that LASeR consistently outperforms the baselines across runs, with standard deviations below 0.3% (lower than RM Agreement Ensemble), showing stable performance. We believe these consistent gains across tasks and models support the claim that LASeR is not only superior but also reliable in expectation.
Table 4: Mean and standard deviation across methods on reasoning tasks.
Method StrategyQA GSM8K MMLU
Best RM 84.26 $\pm$ 0.09 73.10 $\pm$ 0.22 67.11 $\pm$ 0.13
RM Agree. Ensemble 84.05 $\pm$ 0.26 73.65 $\pm$ 0.42 68.07 $\pm$ 0.44
LASeR 85.87 $\pm$ 0.11 74.95 $\pm$ 0.27 68.32 $\pm$ 0.26
LASeR’s Selected RM Adjusts to the Query.
We observe a similar trend in LASeR’s RM selection on LongBench. We observe distinct utilization patterns for the QA tasks vs. summarization and few-shot learning. QA tasks exhibit nearly equal utilization of the top-2 RMs on RewardBench (Zephyr-7B-Alpha and Qwen1.5-7B-Chat in decreasing order), with the utilization of the Qwen RM even exceeds that of Zephyr RM for multi-document QA. In contrast, on summarization and few-shot learning the top RM (Zephyr) is far more preferred by LASeR with margins of $59\%$ and $31\%$ over the second-best RM and the least performant RM being utilized less that $3\%$ of the times.
@@
To further validate our findings, we conducted an additional experiment on RewardBench data. In this setup, we used the MAB trained on Wildchat to select the most appropriate RM for each query. This selected RM was then used to choose the preferred response, which we compared against the ground-truth preferences in RewardBench. We also report the accuracy of each RM in selecting the correct preferred responses. The results are shown in Fig. 8, which demonstrates that LASeR consistently selects the RM whose scoring aligns most closely with the ground-truth preferences.
Figure 8: Accuracy (%) of each RM on selecting the preferred responses on RewardBench. LASeR achieves the highest accuracy.
MAB Rewards Stabilize Over Time.
We observed empirically that the average return (i.e., normalized MAB rewards) of each iteration tends to stabilize over time, particularly after 6-8 training iterations. Fig. 9 shows per-RM normalized MAB reward over training iterations on GSM8K. Notably, the more effective RMs (e.g., Qwen and Zephyr for math reasoning tasks) retain higher MAB rewards over time, suggesting that convergence does not imply uniformity, but rather stabilization of relative effectiveness (this is also consistent with results in Fig. 6 and Fig. 7 that RM selection adapts to the best RM for each task).
Figure 9: Average MAB rewards of different RMs over iterations on GSM-8K.
LASeR Benefits Cold-start Training.
As MAB rewards tend to converge over training iterations, the trained MAB on one dataset can be used for cold-start training on another dataset without retraining the MAB. We conducted an additional experiment where we reinitialized the LLM from the SFT checkpoint and re-trained on reasoning tasks using the converged MAB policy from a prior LASeR run on Wildchat. We denote this new method as cold-start LASeR (CS-LASeR) and present the result in the table below. Results show a comparable performance compared to LASeR (76.32% vs 76.24% on average of 3 tasks), suggesting that using a fixed, learned RM selector policy can benefit cold-start training.
Table 5: Performance of CS-LASeR on reasoning tasks.
Method StrategyQA GSM8K MMLU Avg.
SFT 80.41 69.43 65.66 71.83
LASeR 85.96 74.75 68.24 76.32
CS-LASeR 85.78 74.58 68.37 76.24
Detailed Results for Each RM.
Here, we provide detailed reasoning results for each chosen RM where we use a single RM during training (c.f. Sec. 3.1) in Table 6. These results demonstrate that Qwen1.5-7B-Chat outperforms other RMs on StrategyQA and MMLU, whereas on GSM8K Zephyr-7B-Alpha has the best performance with Llama-3-8B. However, LASeR still yields the best performance, outperforming all RMs by at least $1\%$ on average across reasoning tasks, without the knowledge of which RM is most suited for each task a priori. This is consistent with the results on Wildchat in Table 7 where LASeR outperforms all RMs with significant win rates (e.g., 63.47% and 59.02% against Qwen and Eurus, respectively).